# COPAL: Continual Pruning in Large Language Generative Models

### v1

@@ 1 Introduction
The advent of Large Language Models (LLMs) such as GPT-3 (Brown et al., 2020) and BERT (Devlin et al., 2018) has been a landmark in natural language processing (NLP). Adapting these pre-trained LLMs to diverse domains has offered unprecedented capabilities in various NLP tasks including language understanding and generation (Gururangan et al., 2020). However, there are two primary challenges associated with this approach: (i) extensive computational requirements; and (ii) limited model adaptability. Considering the large-scale models and datasets, the retraining procedure necessitates considerable computational effort, which often limits its feasibility in resource-constrained environments. Moreover, once the models are updated for certain tasks or domains, they may not exhibit the same level of performance when confronted with data that deviate significantly from what they were trained on (known as ”catastrophic forgetting”).
@@
  • Through empirical evaluations on large-scale language models, including LLAMA-7B, 30B, 65B, we show that COPAL outperforms baseline methods, setting a new standard in LLM optimization for both efficiency and adaptability.
@@ 2 Prior Works
Stages of Pruning: Pruning can be applied before training, during training, or post-training. Each stage offers unique advantages and challenges. For instance, SNIP (Lee et al., 2018) and GraSP (Wang et al., 2020) focus on identifying critical connections before training starts. On the other hand, dynamic sparse training (Liu et al., 2020) and soft filter pruning (He et al., 2018) integrate pruning into the training process. Post-training techniques like those developed by Sun et al. (2023) and Frantar & Alistarh (2023) are particularly relevant for large language models, enabling efficient pruning after the model has been fully trained and avoid re-training burden.
@@
Architectural Methods: Adapting the network architecture for continual learning is another approach. Techniques like Progressive Neural Networks (Rusu et al., 2016) and Dynamically Expandable Networks (Yoon et al., 2017) focus on evolving the model’s structure to accommodate new tasks and knowledge.
Continual learning, with its need for finetuning on new datasets, not only risks diminishing the base model’s generalizability but also faces heightened challenges due to rising training costs of LLMs and the evolving nature of real-world NLP applications.
Pruning in Continual Setting. In the field of continual learning, the role of pruning is useful for maintaining and evolving neural network structures effectively. The research by Yan et al. (2022) on Bayesian sparse networks in continual learning settings has been impactful. Similarly, the work of Dekhovich et al. (2023) highlights the importance of preserving efficiency in class-incremental learning. Although these studies emphasize the importance of pruning to keep neural networks efficient and adaptable to new tasks, the potential reduction in a base model’s generalizability and increasing training costs of LLMs training are inevitable. Unlike all existing techniques, we tackle these challenges by proposing a new method for optimizing LLMs, which bypasses the need for re-training as a notion of continual pruning.
@@ 3.2 Weight Stasis in Continual Pruning
To achieve a specific sparsity ratio $s$, the threshold $\mathcal{T}_{s}$ value is found automatically as detailed in Appendix F.
@@
$\begin{split}\mathcal{M}_{i+1}&=\mathcal{I}\left(\mathbf{W}^{*}_{i+1}<\mathcal{T}_{s}\right)\\ &=\mathcal{I}(\left|\mathbf{W}_{i}\cdot\mathcal{R}_{i+1}\right|\cdot\left|\mathcal{M}_{i}\right|<\mathcal{T}_{s})=\mathcal{M}_{i},\\ \mathbf{W}_{i+2}&=\mathbf{W}_{i+1}=\mathbf{W}_{i}\cdot\mathcal{M}_{i}.\\ \end{split}$ (2)
@@
[…]
Since the masked weights in $w^{*}_{i+1}$ are zero for the previously pruned weights ($\mathbf{W}_{i}$), these weights will again fall below the new threshold and remain pruned, regardless of the new dataset $i+1$. Therefore, $\mathbf{W}^{p}_{i+1}$ is being equal to $\mathbf{W}^{p}_{i}$ freezing these weights when transitioning to a new dataset. leading to the phenomenon of ”weight stasis”. The full proof for the inequality is detailed in Appendix G. This limits the model’s ability to adapt to new datasets while adopting calibration-guided pruning strategies to this domain. This is also observed from the experimental results in Table 1.
@@ 4.2 Identification of Crucial Weights
We introduce $\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}$ to capture the sum of the absolute values of the individual gradients until dataset $k$. This metric is particularly insightful for understanding how sensitive the loss is to each individual sample, without regard to the direction of that sensitivity (positive or negative). By focusing on the magnitude rather than the direction, we aim to offer a more nuanced view of the model’s robustness to perturbations in $\mathbf{W}$ and $\mathbf{x}$. The proof is in Appendix D. ~$\tilde{\mathcal{L}}^{k}$ is loss function for only dataset $k$.
$\begin{split}\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}&=\sum_{i=0}^{k}\sum_{j}\left|\nabla_{d\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\nabla^{\prime}_{d\mathbf{W}}\tilde{\mathcal{L}}^{k}+\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k-1}.\end{split}$ (10)
Finally, we use magnitude of directional derivative $D$ of $\mathcal{L}^{i}_{j}$ along $\mathbf{W}$ as our measure of importance weights denoted by $\mathbf{W}_{k}^{*}$, refer to Appendix E for details on directional derivative and the proof.
$\begin{split}\mathbf{W}^{*}_{k}&=\sum_{i=0:k}\sum_{j}\left|D_{\mathbf{W}}\mathcal{L}^{i}_{j}\right|=\left|\mathbf{W}\right|\cdot\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}\\ &=\sum_{j}\left|\mathbf{W}\cdot\nabla_{d\mathbf{W}}\mathcal{L}^{k}_{j}\right|+\mathbf{W}^{*}_{k-1}\end{split}$ (11)
@@ 5.1 Experimental Setting
Models, Datasets. In this paper, we conducted a comprehensive series of experiments utilizing the LLAMA model family, which includes models with 7B, 30B, and 65B parameters, as recently introduced by Touvron et al. (2023). This family of models represents a significant advancement in language model capabilities, offering a range of scales to cater to different computational requirements and performance benchmarks.
Our experimentation focused on three of the most commonly used language datasets in the field: Wikitext-2, the Penn Treebank (PTB), and the Colossal Clean Crawled Corpus (C4). Wikitext-2, introduced by Merity et al. (2016), is renowned for its collection of high-quality, well-structured textual data, predominantly comprising Wikipedia articles. The PTB dataset is a widely-used resource for syntactic analysis (Marcus et al., 1993). Lastly, the C4 dataset, part of the T5 model training corpus, provides a broad and diverse range of internet text, essential for evaluating model performance across various linguistic contexts (Raffel et al., 2020). Our experiments were designed to assess the performance of the language models on these datasets under continual pruning setting.
Baselines. In our comparison, we evaluate the standard magnitude pruning approach, as established by Zhu & Gupta (2017), alongside the more recent developments in post-training pruning works WANDA (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023). Each of these techniques is applied in a layer-wise manner, facilitating scalability even in the context of exceptionally large models.222 More details about the baselines are mentioned in the Appendix I.
@@
LLAMA-7B LLAMA-30B LLAMA-65B
a-bwt m-bwt a-ppl m-ppl a-bwt m-bwt a-ppl m-ppl a-bwt m-bwt a-ppl m-ppl
Dense (no pruning) - - 7.714 10.120 - - 6.131 8.159 - - 6.139 8.878
Unstructured
Magnitude WS WS 30.246 49.670 WS WS 10.958 14.638 WS WS 9.399 13.701
SparseGPT 0.591 0.690 10.166 13.253 0.395 0.730 7.452 9.520 0.334 0.564 7.235 9.874
Wanda 0.569 1.072 9.991 13.626 0.132 0.192 7.261 9.231 0.172 0.628 7.584 11.354
COPAL (ours) 0.016 0.032 9.728 12.585 0.007 0.025 7.240 9.081 0.001 0.208 6.791 8.839
Semi structured 2:4
Magnitude WS WS 131.653 303.710 WS WS 13.757 19.139 WS WS 10.544 14.704
SparseGPT 4.365 6.391 15.744 21.771 1.436 3.015 9.592 12.657 2.128 6.979 9.333 16.037
Wanda 1.667 3.192 16.154 23.266 0.493 1.079 9.363 12.183 0.184 0.389 9.230 12.665
COPAL (ours) 0.009 0.075 15.335 21.159 0.036 0.100 9.274 11.478 0.038 0.258 9.161 11.233
Semi structured 4:8
Magnitude WS WS 32.105 56.652 WS WS 12.998 16.881 WS WS 9.247 12.568
SparseGPT 1.929 3.045 11.924 15.884 0.838 1.670 8.351 10.790 0.907 2.992 8.363 13.358
Wanda 0.771 1.645 11.929 16.631 0.231 0.486 8.094 10.353 0.172 0.610 8.315 11.746
COPAL (ours) 0.038 0.075 11.734 15.532 0.012 0.050 8.032 9.982 0.019 0.260 8.291 10.807
Table 3: Results of per dataset performance continual pruning on wikitext2, ptb, c4 datasets with all incremental permutations on LLAMA-7B,30B,65B. Unstructured pruning with $50\%$ sparsity ratio. Overall Standard deviations in Appendix H
Wikitext2 PTB C4
BWT PPL BWT PPL BWT PPL
LLAMA-7B
Dense (no pruning) - 5.677 ± 0.00 - 10.120 ± 0.00 - 7.344 ± 0.00
Magnitude WS 17.288 ± 0.000 WS 49.671 ± 0.000 WS 23.778 ± 0.000
LLAMA-30B
Dense (no pruning) - 4.101 ± $0$ 8.159 ± $0$ 6.131 ± 0.000
Magnitude WS 7.542 ± 0.000 WS 14.638 ± 0.000 WS 10.696 ± 0.000
LLAMA-65B
Dense (no pruning) - 3.562 ± 0.00 - 8.878 ± 0.00 - 5.978 ± 0.00
Magnitude WS 5.901 ± 0.000 WS 13.701 ± 0.000 WS 8.595 ± 0.000
@@ 5.2 Results
Unstructured Continual Pruning. In Table 3, the LLAMA-7B model under unstructured pruning, COPAL exhibits a pronounced superiority. It drastically lowers average BWT to 0.016, compared to best performing baseline Wanda’s 0.569, demonstrating a 97.2% improvement. This indicates a highly effective pruning process with minimal knowledge loss. In Table 4, COPAL’s PPL performance on individual datasets like Wikitext2 and PTB shows PPL performance improvement of 7.144 (compared to best performing baseline Wanda’s 7.208) and 12.567 (compared to best performing baseline SparseGPT’s 13.141), respectively at 50% sparsity ratio and still shows comparable performance to the base model. This consistency across diverse datasets underlines the method’s robustness and adaptability.
@@
Semi-structured N:M Continual Pruning. In the realm of semi-structured N:M continual pruning, the COPAL technique stands out for its effectiveness across various pruning configurations, notably within the context of the LLAMA-65B model. When applied to a 2:4 pruning pattern, COPAL achieves a 98.2% improvement in average Backward Weight Transfer (BWT), reducing it to 0.038 compared to SparseGPT’s 2.128. Additionally, it enhances the average Perplexity (PPL) by 1.8%, bringing it down to 9.161 from SparseGPT’s 9.333. The proficiency of COPAL is further evidenced in the 4:8 pruning pattern, where it markedly reduces average BWT by 99.3% to 0.019, compared to SparseGPT’s 0.907, and achieves a 0.9% improvement in average PPL, lowering it to 8.291 from 8.363. These outcomes underscore COPAL’s robust adaptability and efficiency in managing semi-structured pruning, particularly within large-scale models like LLAMA-65B, where balancing efficiency and performance is paramount.

### v2

@@ 1 Introduction
The advent of Large Language Models (LLMs) such as GPT-3 (Brown et al., 2020) and LLaMA (Touvron et al., 2023) has been a landmark in natural language processing (NLP). Adapting these pre-trained LLMs to diverse domains has offered unprecedented capabilities in various NLP tasks including language understanding and generation (Gururangan et al., 2020). However, there are two primary challenges associated with this approach: (i) extensive computational requirements; and (ii) limited model adaptability. Considering the large-scale models and datasets, the retraining procedure necessitates considerable computational effort, which often limits its feasibility in resource-constrained environments. Moreover, once the models are updated for certain tasks or domains, they may not exhibit the same level of performance when confronted with data that deviate significantly from what they were trained on (known as ”catastrophic forgetting”).
@@
  • Through empirical evaluations on large-scale language models, including LLaMA-7B, 13B, 30B, 65B, we show that COPAL outperforms baseline methods, setting a new standard in LLM optimization for both efficiency and adaptability.
[…]
@@ 2 Prior Works
Stages of Pruning: Pruning can be applied before training, during training, or post-training. Each stage offers unique advantages and challenges. For instance, SNIP (Lee et al., 2018) and GraSP (Wang et al., 2020) focus on identifying critical connections before training starts. On the other hand, dynamic sparse training (Liu et al., 2020), soft filter pruning (He et al., 2018) or with knowledge distillation(xia2022structured; yang2022gradient) integrate pruning into the training process. Post-training techniques like those developed by Sun et al. (2023) and Frantar & Alistarh (2023) are particularly relevant for large language models, enabling efficient pruning after the model has been fully trained and avoid re-training burden.
@@
Architectural Methods: Adapting the network architecture for continual learning is another approach. Techniques like Progressive Neural Networks (Rusu et al., 2016) and Dynamically Expandable Networks (Yoon et al., 2017) focus on evolving the model’s structure to accommodate new tasks. Continual learning, with its need for finetuning on new datasets, not only risks diminishing the base model’s generalizability but also faces heightened challenges due to rising training costs of LLMs and the evolving nature of real-world NLP applications.
Pruning in Continual Setting. In the field of continual learning, the role of pruning is useful for maintaining and evolving neural network structures effectively. The research by Yan et al. (2022) on Bayesian sparse networks in continual learning settings has been impactful. Similarly, the work of Dekhovich et al. (2023) highlights the importance of preserving efficiency in class-incremental learning. Although these studies emphasize the importance of pruning to keep neural networks efficient and adaptable to new tasks, the potential reduction in a base model’s generalizability and increasing training costs of LLMs training are inevitable. Unlike all existing techniques, we tackle these challenges with a new method for optimizing LLMs, which bypasses the need for re-training as a notion of continual pruning.
@@ 3.2 Weight Stasis in Continual Pruning
To achieve a specific sparsity ratio $s$, the threshold $\mathcal{T}_{s}$ value is found automatically as detailed in Appendix D.
@@
$\begin{split}\mathcal{M}_{i+1}&=\mathcal{I}\left(\mathbf{W}^{*}_{i+1}<\mathcal{T}_{s}\right)\\ &=\mathcal{I}(\left|\mathbf{W}_{i+1}\cdot\mathcal{R}_{i+1}\right|<\mathcal{T}_{s})\\ &=\mathcal{I}(\left|(\mathbf{W}_{i}\cdot\mathcal{M}_{i})\cdot\mathcal{R}_{i+1}\right|<\mathcal{T}_{s})\\ &=\mathcal{I}(\left|\mathbf{W}_{i}\cdot\mathcal{R}_{i+1}\right|\cdot\left|\mathcal{M}_{i}\right|<\mathcal{T}_{s})\\ &\begin{cases}0&\text{if }m<\mathcal{T}_{s},m\in\left|\mathbf{W}_{i}\cdot\mathcal{R}_{i+1}\right|\cdot\left|\mathcal{M}_{i}\right|\\ 1&\text{otherwise.}\end{cases},\\ &=\mathcal{M}_{i},\\ \mathbf{W}_{i+2}&=\mathbf{W}_{i+1}=\mathbf{W}_{i}\cdot\mathcal{M}_{i}.\\ \end{split}$ (2)
@@
Since the masked weights in $w^{*}_{i+1}$ are zero for the previously pruned weights ($\mathbf{W}_{i}$), these weights will again fall below the new threshold and remain pruned, regardless of the new dataset $i+1$. Therefore, $\mathbf{W}^{p}_{i+1}$ is being equal to $\mathbf{W}^{p}_{i}$ freezing these weights when transitioning to a new dataset. leading to the phenomenon of ”weight stasis”. This limits the model’s ability to adapt to new datasets while adopting calibration-guided pruning strategies to this domain. This is also observed from the experimental results in Table 1.
@@ 4.2 Identification of Crucial Weights
We further introduce $\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}$ to capture the sum of the absolute values of the individual gradients until dataset $k$.
$\begin{split}\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}&=\sum_{i=0}^{k}\sum_{j}\left|\nabla_{d\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\sum_{j}2\left|d\mathbf{y}^{k}_{j}\frac{\partial f}{\partial\mathbf{W}}\right|+\sum_{i=0:k-1}\sum_{j}2\left|d\mathbf{y}^{i}_{j}\frac{\partial f}{\partial\mathbf{W}}\right|\\ &=\sum_{j}2\left|d\mathbf{y}^{k}_{j}\right|\left|\frac{\partial f}{\partial\mathbf{W}}\right|+\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}_{k-1}\\ &=\nabla^{\prime}_{d\mathbf{W}}\tilde{\mathcal{L}}^{k}+\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k-1},\end{split}$ (10)
where ~$\tilde{\mathcal{L}}^{k}$ is loss function for only dataset $k$. This metric is particularly insightful for understanding how sensitive the loss is to each individual sample, without regard to the direction of that sensitivity (positive or negative). By focusing on the magnitude rather than the direction, we aim to offer a more nuanced view of the model’s robustness to perturbations in $\mathbf{W}$ and $\mathbf{x}$.
Finally, we use magnitude of directional derivative $D$ of $\mathcal{L}^{i}_{j}$ along $\mathbf{W}$ as our measure of importance weights denoted by $\mathbf{W}_{k}^{*}$.
$\begin{split}\mathbf{W}^{*}_{k}&=\sum_{i=0:k}\sum_{j}\left|D_{\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\sum_{i=0:k}\sum_{j}\left|\mathbf{W}\cdot\nabla_{d\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\left|\mathbf{W}\right|\cdot\sum_{i=0:k}\sum_{j}\left|\nabla_{d\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\left|\mathbf{W}\right|\cdot\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}\\ &=\left|\mathbf{W}\cdot\right|(\nabla_{d\mathbf{W}}\tilde{\mathcal{L}}^{k}+\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}_{k-1})\\ &=\sum_{j}\left|\mathbf{W}\cdot\nabla_{d\mathbf{W}}\mathcal{L}^{k}_{j}\right|+\mathbf{W}^{*}_{k-1}.\end{split}$ (11)
@@ 5.1 Experimental Setting
Models, Datasets. In this paper, we conducted a comprehensive series of experiments utilizing the LLaMA model family, which includes models with 7B, 13B, 30B, and 65B parameters, as recently introduced by Touvron et al. (2023). This family of models represents a significant advancement in language model capabilities, offering a range of scales to cater to different computational requirements and performance benchmarks.
Our experimentation focused on three of the most commonly used language datasets in the field: Wikitext-2 (Merity et al., 2016), the Penn Treebank (PTB) (Marcus et al., 1993), and the Colossal Clean Crawled Corpus (C4) (Raffel et al., 2020). Wikitext-2, introduced by Merity et al. (2016), is renowned for its collection of high-quality, well-structured textual data, predominantly comprising Wikipedia articles. The PTB dataset is a widely-used resource for syntactic analysis (Marcus et al., 1993). Lastly, the C4 dataset, part of the T5 model training corpus, provides a broad and diverse range of internet text, essential for evaluating model performance across various linguistic contexts (Raffel et al., 2020).
Baselines. In our comparison, we evaluate the standard magnitude pruning approach, as established by Zhu & Gupta (2017), alongside the more recent developments in post-training pruning works WANDA (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023). Each of these techniques is applied in a layer-wise manner, facilitating scalability even in the context of exceptionally large models. Magnitude pruning, a method detailed by Zhu & Gupta (2017), effectively compresses models by removing weights with the smallest absolute values, which are deemed least impactful on the network’s output. In contrast, SparseGPT, introduced by Frantar & Alistarh (2023), integrates sparsity into the post-training process of transformer-based models. It often employs methods such as utilizing the Hessian matrix to identify weights that can be pruned with minimal loss in performance, effectively making the model sparse from the outset. Finally, WANDA, as proposed by Sun et al. (2023), does pruning by analyzing weight and activation distributions. This technique identifies and prunes network segments that minimally contribute to outputs.
In Table 2, baseline models employ global initialization due to weight stasis in sequential setups, as detailed in Section 3.2 and Table 1. Global initialization also leads to better average perplexity compared to sequential initialization.
@@
LLAMA-7B LLAMA-30B
a-bwt m-bwt a-ppl m-ppl a-bwt m-bwt a-ppl m-ppl
Dense (no pruning) - - 7.714 10.120 - - 6.131 8.159
Unstructured
Magnitude WS WS 30.246 49.670 WS WS 10.958 14.638
SparseGPT 0.591 0.690 10.166 13.253 0.395 0.730 7.452 9.520
Wanda 0.569 1.072 9.991 13.626 0.132 0.192 7.261 9.231
COPAL (ours) 0.016 0.032 9.728 12.585 0.007 0.025 7.240 9.081
Semi structured 2:4
Magnitude WS WS 131.653 303.710 WS WS 13.757 19.139
SparseGPT 4.365 6.391 15.744 21.771 1.436 3.015 9.592 12.657
Wanda 1.667 3.192 16.154 23.266 0.493 1.079 9.363 12.183
COPAL (ours) 0.009 0.075 15.335 21.159 0.036 0.100 9.274 11.478
Semi structured 4:8
Magnitude WS WS 32.105 56.652 WS WS 12.998 16.881
SparseGPT 1.929 3.045 11.924 15.884 0.838 1.670 8.351 10.790
Wanda 0.771 1.645 11.929 16.631 0.231 0.486 8.094 10.353
COPAL (ours) 0.038 0.075 11.734 15.532 0.012 0.050 8.032 9.982
LLAMA-13B LLAMA-65B
a-bwt m-bwt a-ppl m-ppl a-bwt m-bwt a-ppl m-ppl
Dense (no pruning) - - 6.990 9.081 - - 6.139 8.878
Unstructured
Magnitude WS WS 28.935 41.368 WS WS 9.399 13.701
SparseGPT 0.606 1.123 8.467 11.094 0.334 0.564 7.235 9.874
Wanda 0.203 0.298 8.570 10.896 0.172 0.628 7.584 11.354
COPAL (ours) 0.029 0.078 8.354 10.818 0.001 0.208 6.791 8.839
Semi-structured (2:4)
Magnitude WS WS 28.702 44.072 WS WS 10.544 14.704
SparseGPT 2.670 5.978 11.970 17.386 2.128 6.979 9.333 16.037
Wanda 0.871 1.698 13.209 18.920 0.184 0.389 9.230 12.665
COPAL (ours) -0.007 0.340 11.455 17.155 0.038 0.258 9.161 11.233
Semi-structured (4:8)
Magnitude WS WS 20.476 29.055 WS WS 9.247 12.568
SparseGPT 1.248 2.905 9.773 13.456 0.907 2.992 8.363 13.358
Wanda 0.375 0.793 9.946 13.281 0.172 0.610 8.315 11.746
COPAL (ours) -0.019 0.160 9.402 12.338 0.019 0.260 8.291 10.807
Table 3: Results of per dataset performance continual pruning on wikitext2, ptb, c4 datasets with all incremental permutations on LLaMA-7B,13B,30B,65B. Unstructured pruning with $50\%$ sparsity ratio. Overall Standard deviations in Appendix E
Wikitext2 PTB C4
BWT PPL BWT PPL BWT PPL
LLAMA-7B
Dense (no pruning) - 5.677 ± $0$ 10.120 ± $0$ 7.344 ± 0.000
Magnitude WS 17.288 ± 0.000 WS 49.671 ± 0.000 WS 23.778 ± 0.000
LLAMA-13B
Dense (no pruning) - 5.091 ± $0$ 9.081 ± $0$ 6.798 ± 0.000
Magnitude WS 20.215 ± 0.000 WS 41.368 ± 0.000 WS $2$ 0.000
SparseGPT 0.465 ± 0.093 6.301 ± 0.241 1.121 ± 0.002 10.532 ± 0.561 0.233 ± 0.049 8.568 ± 0.121
Wanda 0.165 ± 0.040 6.138 ± 0.087 0.285 ± 0.013 10.663 ± 0.143 0.159 ± 0.028 8.262 ± 0.082
COPAL (ours) 0.052 ± 0.022 6.096 ± 0.032 0.011 ± 0.018 10.472 ± 0.020 0.023 ± 0.026 8.176 ± 0.019
LLAMA-30B
Dense (no pruning) - 4.101 ± $0$ 8.159 ± $0$ 6.131 ± 0.000
Magnitude WS 7.542 ± 0.000 WS 14.638 ± 0.000 WS 10.696 ± 0.000
LLAMA-65B
Dense (no pruning) - 3.562 ± $0$ 8.878 ± $0$ 5.978 ± 0.000
Magnitude WS 5.901 ± 0.000 WS 13.701 ± 0.000 WS 8.595 ± 0.000
@@ 5.2 Results
Unstructured Continual Pruning. In Table 2, the LLaMA-7B model under unstructured pruning, COPAL exhibits a pronounced superiority. It drastically lowers average BWT to 0.016, compared to best performing baseline Wanda’s 0.569, demonstrating a 97.2% improvement. This indicates a highly effective pruning process with minimal knowledge loss. In Table 3, COPAL’s PPL performance on individual datasets like Wikitext2 and PTB shows PPL performance improvement of 7.144 (compared to best performing baseline Wanda’s 7.208) and 12.567 (compared to best performing baseline SparseGPT’s 13.141), respectively at 50% sparsity ratio and still shows comparable performance to the base model. For the LLaMA-13B model, COPAL reduces average BWT to 0.029 and average PPL to 8.354, indicating substantial improvements in efficiency and comprehension. This consistency across diverse datasets underlines the method’s robustness and adaptability.
@@
Semi-structured N:M Continual Pruning. In the realm of semi-structured N:M continual pruning, the COPAL technique stands out for its effectiveness across various pruning configurations, notably within the context of the LLaMA-65B model. When applied to a 2:4 pruning pattern, COPAL achieves a 98.2% improvement in average Backward Weight Transfer (BWT), reducing it to 0.038 compared to SparseGPT’s 2.128. Additionally, it enhances the average Perplexity (PPL) by 1.8%, bringing it down to 9.161 from SparseGPT’s 9.333. The proficiency of COPAL is further evidenced in the 4:8 pruning pattern, where it markedly reduces average BWT by 99.3% to 0.019, compared to SparseGPT’s 0.907, and achieves a 0.9% improvement in average PPL, lowering it to 8.291 from 8.363. In the 2:4 pruning configuration for LLaMA-13B, COPAL achieves BWT of -0.007 and PPL of 11.455, illustrating a marked enhancement in the model’s pruning efficacy. These outcomes underscore COPAL’s robust adaptability and efficiency in managing semi-structured pruning, particularly within large-scale models like LLaMA-65B, where balancing efficiency and performance is paramount.